Parth (parthkohli):

Nice little problem. Show that \(\dfrac{(ax - b)(dx -c)}{(bx - a)(cx - d)}\) will be capable of all real values if \(a^2 - b^2\) and \(c^2 - d^2\) have the same sign.

Parth (parthkohli):

@mathslover and @mukushla try this.

OpenStudy (anonymous):

capable ??

Parth (parthkohli):

@amistre64

Parth (parthkohli):

Can I give the solution off now? Or is it too early?

OpenStudy (amistre64):

if you want to give the solution, especially to your own question, then feel free :)

OpenStudy (amistre64):

if you want it to sit for a few days, bumping, then thats fine too

Parth (parthkohli):

Hehe, I'm too impatient. :P Any ideas about the problem?

OpenStudy (anonymous):

i would try \(\iff\) i know what do you mean by capable

OpenStudy (experimentx):

for what domain of x?

Parth (parthkohli):

Real.

OpenStudy (experimentx):

you are essentially saying that this maps whole R to R.

Parth (parthkohli):

@BSwan It means that this function can return any possible real number for a particular \(x\) in reals.

OpenStudy (experimentx):

well well ... let y=that function and solve the quadratic equation.

Parth (parthkohli):

Haha nice.

OpenStudy (experimentx):

i can see that there exists a map from R to R, but I am not sure if it's one to one map.

OpenStudy (experimentx):

it's because there exists a thing called pole of order 1.

OpenStudy (experimentx):

actually two poles that maps to +infinity and -infinity ,,, i am guessing this is not one to one function.

Parth (parthkohli):

\[f(1) = 1\]\[f(-1)=1\]

OpenStudy (experimentx):

interesting

Parth (parthkohli):

so it's not a one-to-one function. :)

OpenStudy (experimentx):

yep!!

Parth (parthkohli):

anyway, experimentX's solution is right. 1. Suppose that the function is \(y\). 2. Take the denominator to the other side. 3. Rearrange and obtain an expression in the form \(ax^2 + bx + c = 0\). \(a,b,c\) can contain terms in \(y\). 4. You basically have to find the condition for which this quadratic has real roots. Use the discriminant. This will leave you with an expression that is \(\ge 0\). 5. Manipulate that expression until you get \((a^2 - b^2)(c^2 - d^2) \ge 0\).

OpenStudy (experimentx):

well well ... nice job there (y) I suppose I am just too lazy :(